The number of visits to a website in one hour has the following probability mass function


{1/10 , π‘₯ = 0

𝑝(x)= { π‘˜π‘₯/10 , π‘₯ = 1,2,3

{ π‘˜(6βˆ’π‘₯)/15 , π‘₯ = 4,5



a) Determine the value of π‘˜.


With the value of π‘˜ obtained:


b) Find the probability of having

i) 3 visits in one hour.

ii) less than 2 visits in one hour.

iii) more than 6 visits in two hours.

c) By using central limit theorem, estimate the probability of having less than 253 visits in 100 hours.


1
Expert's answer
2020-08-14T17:34:27-0400

Solution:

a)

The sum of all probability is 1.

So,

P(X=0)+P(X=1)+P(X=2)+P(X=3)+P(X=4)+P(X=5)=1.

1/10+k/10+2k/10+3k/10+2k/15+k/15=1.

1/10+3k/5+k/5=1

4k/5=9/10

k=9/8.

So,

P(X=0)=1/10,

P(X=1)=9/80,

P(X=2)=9/40,

P(X=3)=27/80,

P(X=4)=3/20,

P(X=5)=3/40.

b)

i) P(X=3)=27/80,

The probability of Β 3 visits in one hour is equal to 27/80.

ii) P(X<2)=P(X=0)+P(X=1)=1/10+9/80=17/80,

The probability of Β event that less than 2 visits in one hour is equal to 17/80.

iii) k1+k2>6:

(k1,k2)=(2,5),(3,4),(3,5),(4,3)(4,4),(4,5),(5,5)

P(X1+X2>6)=P(X1=2,X2=5)+P(X1=3,X2=4)+P(X1=3,X2=5)+P(X1=4,X2=3)+P(X1=4,X2=4)+P(X1=4,X2=5)+P(X1=5,X2=5)

P(X1+X2>6)=(9/40)(3/40)+(27/80)(3/20)+(27/80)(3/40)+(3/20)(27/80)+(3/20)(3/20)+(3/20)(3/40)+(3/40)(3/40)=117/640

The probability of Β event that more than 6 visits in two housr is equal to 117/640.


c) "\mu=0\cdot P(X=0)+1\cdot P(X=1)+2\cdot P(X=2)+3\cdot P(X=3)+4\cdot P(X=4)+5\cdot P(X=5)=0\cdot 1/10+1\cdot 9/80+2\cdot 9/40+3\cdot 27/80+4\cdot 3/20+5\cdot 3/40=51/20."

"\mu=51/20."

"\sigma^2=E[X^2]-\mu^2"

"E[X^2]=0\cdot P(X=0)+1\cdot P(X=1)+4\cdot P(X=2)+9\cdot P(X=3)+16\cdot P(X=4)+25\cdot P(X=5)=0\cdot 1/10+1\cdot 9/80+4\cdot 9/40+9\cdot 27/80+16\cdot 3/20+25\cdot 3/40=333/40."

"\sigma^2=333/40-(51/20)^2=729/400"

"\sigma=\sqrt{729/400}=27/20."

"n=100"

"\beta=(253-n\cdot\mu)/(\sigma\cdot n^{1/2})"

"\beta=(253-100\cdot\mu)/(10\sigma)=-4/27"

"P(S_{n}<253)=P((S_{n}-n\cdot\mu)/(\sigma\cdot n^{1/2})<(253-n\cdot\mu)/(\sigma\cdot n^{1/2}))"


"P(S_{n}<253)=P((S_{n}-n\cdot\mu)/(\sigma\cdot n^{1/2})<-4/27)=\Phi(-4/27)"

"P(S_{100}<253)=\Phi(-4/27)"

"\Phi(-4/27)\approx0.4411" (Here we use Wolfram Alpha)


The probability of having less than 253 visits in 100 hours is approximately equal to 0.4411.